### v1

# Nonlinear integral extension of PID control with improved convergence of perturbed second-order dynamic systems

Abstract
Nonlinear extension of the integral part of PID feedback control is proposed for the perturbed second-order systems. For the matched constant perturbations, the global asymptotic stability is shown, and for Lipschitz perturbations an ultimately bounded output error is guaranteed. The second-order system plants can also be expanded by an additional (parasitic) actuator dynamics with low-pass characteristics. The proposed nonlinear control is proven to outperform its linear (PID) benchmarking counterpart during the settling phase, i.e. at convergence of the residual output error. An experimental case study of the second-order system with an additional actuator dynamics and considerable perturbation is demonstrated to confirm and benchmark the control performance.
I INTRODUCTION
Proportional-integral-derivative (PID) controllers, see e.g. [1, 2], as well as their possible extensions like anti-windup, gain-scheduling, parameters adaptation, are still mostly seen as a working horse in various industries. Despite a large number of available and well perceived tutorials and surveys on PID principles, tuning, extensions, and limitations, e.g. [3, 4, 5] just for mentioning here the few, numerous applications face (here and again) some crucial challenges with use of the PID controllers. These lie often in a trade-off between the performance and robustness, while the metrics associated with a transient and steady-state response appear to be most significant for several applications.
Well known, an integral control action is required for asymptotic set-point regulation under parameter perturbations or (equally usual in practice) presence of the matched stationary disturbances. For simple first- and second-order processes, including possibly also time delays, the tuning of the integral time constant is straightforward and well understood, see e.g. [6]. More challenging becomes an effective integral control action in case of disturbance processes that vary over time, especially in vicinity of the set-point, and when the application requires an increased settling accuracy in shorter time. Not surprising is, hence, that a standard integral control action cannot, for example, compensate accurately for the nonlinear Coulomb friction [7, 8] in the set-point tasks. As one of the possible solutions, the so-called reset integral control (introduced originally in [9]) can be applied, see e.g. [10]. The reset integral behavior in feedback, and thus a switching dynamics, restricts however the applicability to a certain class of the system plants [11]. Worth noting is that despite a relaxation applied to the underlying base systems and, this way, an attempt to address the stability by a less stringent circle-criterion conditions [12], the reset integrator does not entirely satisfy the necessary conditions posed on the feedback nonlinearity, cf. [13, 14], cf. with preliminaries provided below in section II. Another approach of the time-varying integral control gain, which assumes a monotonic decrease of the gain to zero at a sufficiently slow rate, was presented in [15] for linear system plants subject to a globally Lipschitz and non-decreasing actuator nonlinearity. Despite the profound theoretical results, which also form basis for a simple adaptive control strategy, the approach is not directly applicable to compensate for external disturbances and to improve the convergence performance. An universal integral controller for minimum-phase nonlinear systems and the tuning procedures for its key parameters were proposed in [16]. A single-input-single-output (SISO) nonlinear system plant that has a well-defined normal form with asymptotically stable zero dynamics was assumed for the synthesis procedure. Remarkable is that for relative degree-one systems, this universal controller reduces to the classical PI controller followed by saturation, and for relative degree-two systems it reduces to the classical PID controller followed by saturation. Settling performance, in terms of the convergence time and accuracy were, however, not explicitly in focus of the proposed methodology. Later, the problem of robust nonlinear integral control for a class of systems with dynamic uncertainties was addressed in [17], while a constructive controller design procedure was presented in the presence of an unknown equilibrium due to uncertain nonlinearities and dynamic uncertainties. The resulted controller reduces, notwithstanding, to linear control laws if the nominal system is linear, and becomes a traditional PI control if, additionally, the relative degree is one, or the uncertainty falls into the input space, i.e. strict matching condition.
Motivated by the background provided above, and the steady-state behavior of motion control systems, see e.g. [18], the present work introduces a possible nonlinear integral extension of the PID feedback control. To the best of the author’s knowledge, the proposed control, see eq. (8), is new and was not used in the previously published works.
The rest of the paper is structured in the following way. Section II provides the notation in use and the necessary preliminaries of a standard PID feedback control and circle criterion of absolute stability. The main results are placed into section III, starting with absolute stability, followed by the disturbance rejection and extension of the second-order system plant by an additional actuator dynamics. Also some clarifying numerical examples are included. An experimental case study used for evaluation of the proposed control and benchmarking it with the standard PD and PID controllers is provided in section IV. The paper ends with the short conclusions drawn in section V.
II PRELIMINARIES
Notations
Unless otherwise said, $x=[x_{1},x_{2},\ldots,x_{n}]^{\top}\in\mathbb{R}^{n}$ is a state vector of real numbers, $t\in\mathbb{R}_{+}=\{\tau\in\mathbb{R}\>:\>\tau\geq 0\}$ is a non-negative time variable. The time derivative of a variable $x_{n}$ is denoted by $\dot{x}_{n}\equiv dx_{n}/dt$, respectively the second time derivative is denoted by $\ddot{x}_{n}\equiv d^{2}x_{n}/dt^{2}$, and so on. An input-output transfer function is denoted by $H(s)=y(s)/u(s)$, where $s=(\rho+j\omega)\in\mathbb{C}$ is the complex Laplace variable with the signed angular frequency $\omega\in\mathbb{R}$, imaginary unit $j^{2}=-1$, and $\rho\in\mathbb{R}$. An absolute value in $\mathbb{R}$ is denoted by ||$\cdot$|| ⋅ |, while $\bigl{|}H(s)\bigr{|}$ and respectively $\bigl{|}H(j\omega)\bigr{|}$ denote the absolute value of a complex function. An identity matrix of the appropriate dimension is denoted by $I$.
II-A PID controlled feedback loop
We consider the second-order systems of the type
$\ddot{y}(t)=u(t)+\sigma(t),$ (1)
where the output of interest $y(t)$ and its time derivative $\dot{y}(t)$ are available for a feedback control. The control signal $u(t)\in\mathbb{R}$ is continuous on $t\in\mathbb{R}_{+}$ and assumed to stay unsaturated [^1] We notice that a feedback control involving an integral action may be subject to saturation, which is also quite common practice in control engineering. However, an explicit measure against the associated wind-up effects is beyond the scope of the current work. Therefore, neither standard PID control, assumed as a reference benchmarking control system, nor its nonlinear extension proposed handle a saturated operation mode. for $(y,\dot{y})\in\mathbb{R}^{2}$. The matched perturbation function $\sigma(t)$ is assumed to be Lipschitz, i.e. $|\dot{\sigma}|<\Sigma$.
First, we recall briefly a standard PID feedback control
[…]
with the design parameters $a,b,c>0$, while focusing below on the most relevant (for us) properties and features. Control parameters are assumed to be appropriately determined.
One can easily show, either in time or frequency domain, that for a constant disturbance, i.e. $\dot{\sigma}=0$, the PID feedback control (2) provides a global asymptotic stabilization of (1). For a Lipschitz disturbance, mapped into frequency domain by $\Sigma\propto\omega$, the PID feedback control (2) provides an ultimately bounded output error $|y(t)|<\varepsilon(\omega)$ for $t\rightarrow\infty$. Regarding this statement, which is rather self-understood in the linear feedback control theory, cf. e.g. [19, 2], we will have a closer look below in section III-B.
Substituting (2) into (1) and analyzing the characteristic polynomial of the unperturbed system dynamics, i.e.
$s^{3}+as^{2}+bs+c=0,$ (3)
one can show by, for example Routh-Hurwitz stability criterion, see e.g. [19], that the closed-loop control system is asymptotically (and exponentially) stable if and only if
$ab-c>0.$ (4)
That means every solution of the state trajectories is uniformly bounded and satisfies $y(t)\rightarrow 0$, $\dot{y}(t)\rightarrow 0$ as $t\rightarrow\infty$.
II-B Circle criterion
For a linear time invariant (LTI) system in the minimal state-space realization, see e.g. [20] for detail, the nonlinearity $\phi(\cdot,\cdot)$ that is acting in feedback and satisfying
$\phi(0,t)$ $=$ $0\quad\forall\;t\in\mathbb{R}_{+},$
$z\,\phi(z,t)$$\geq$ $0\quad\forall\;z\in\mathbb{R},\>t\in\mathbb{R}_{+}$
results in the so called Lur’e problem [21]. The latter is also refereed to as the absolute stability problem and is formulated for a generic class of systems given by
$\dot{x}(t)$ $=$ $Ax(t)+Bu(t),$
$z(t)$ $=$ $Cx(t)+Du(t),$ (5)
$u(t)$ $=$ $-\phi\bigl{(}z(t),t\bigr{)}.$
Note that the single-input-single-output (SISO) systems are considered, so that $u,\,z\in\mathbb{R}$, while $A,\,B,\,C,\,D$ are the matrices (correspondingly vectors) of an appropriate dimension and contain the system parameters. Here and for the rest of the section, we will closely follow the notations and developments of circle criterion [22, 23] as these are provided in [14]; for additional explanations see also [24].
For applying the circle criterion of absolute stability to (II-B), the following necessary assumptions are placed.
  1. i)
    $A$, $B$, and $C$ are the minimal realization of a linear subsystem in (II-B), i.e. the pair $(A,B)$ is fully controllable and the pair $(A,C)$ is fully observable.
  2. ii)
    The system matrix $A$ has all eigenvalues $\lambda_{i}$ lying in $\mathbb{C}_{-}^{\circ}$, i.e. $\mathrm{Re}\bigl{[}\lambda_{i}(A)\bigr{]}<0$ for $i=1,\ldots,n$.
  3. iii)
    The nonlinearity $\phi(\cdot,t)$ belongs to the sector $[k_{1},k_{2}]$ for $k_{2}\geq k_{1}\geq 0$. That means
    $k_{1}z\leq\phi(z,t)\leq k_{2}z.$
Moreover, it is assumed that
  1. iv)
    $\phi(z,t)$ is memoryless, possibly time-varying, and piecewise continuous in $t$ and locally Lipschitz in $z$.
For the input-output transfer function
$H(s)=C(sI-A)^{-1}+D$ (6)
of the linear subsystem in (II-B), the absolute stability of the closed-loop (II-B) with $\phi(\cdot,t)$ belonging to the sector $[0,k]$, is guaranteed by the circle criterion stated in the below theorem.
Theorem 1
Given a feedback system described by the equations (II-B). Let the assumptions (i)-(iv) be effective. Then, the origin of (II-B) is globally asymptotically stable if
$\mathrm{Re}\bigl{[}1+kH(j\omega)\bigr{]}>0,\quad\forall\>\omega.$ (7)
The proof of Theorem 1 can be found in [14, page 241].
It is worth recalling that the circle criterion given by (7) for the sector $[0,k]$ is equivalent to, correspondingly can be transformed into (and also from), the circle criterion formulated for the sector $[k_{1},k_{2}]$ in terms of a disc $\mathcal{D}(k_{1},k_{2})$ in the polar coordinates of $H(j\omega)$-locus, see e.g. [24, 14].
III MAIN RESULTS
The proposed feedback control law with nonlinear extension of the integral part is given by, cf. (2),
[…]
The linear control parameters $a,b,c>0$ are in the same sense as for (2). The nonlinear parameters $d\geq 0$ and $e<0$ provide a smooth variation of the overall integral gain
$\Omega\equiv c\,\Bigl{(}1+d\exp\bigl{(}e|y(t)|\bigr{)}\Bigr{)},$ (9)
that is monotonically increasing with decrease of the magnitude of the control error. Worth emphasizing is that
  1. (i)
    $c\leq\Omega\leq c+cd$;
  2. (ii)
    $\Omega\rightarrow c$ for a continuously growing $|y|$;
  3. (iii)
    the control (8) recovers completely to the benchmarking reference controller (2) if $d=0$.
In the following, we are proving the absolute stability of an unperturbed system (1), (8) by using the circle criterion, cf. with section II-B. Afterwards, the closed-loop system (1), (8) with an exogenous input $\sigma$ is analyzed in terms of the convergence performance of $y(t)$, and that in comparison to the control law (2), i.e. for $d=0$ in (8).
III-A Absolute stability
The unperturbed closed-loop system (1), (8) is transferable into the Lur’e problem (II-B) with $D=0$ and
$A=\left(\begin{array}[]{ccc}0&1&0\\ 0&0&1\\ -c&-b&-a\\ \end{array}\right),\,B=\left(\begin{array}[]{c}0\\ 0\\ 1\\ \end{array}\right),\,C^{\top}=\left(\begin{array}[]{c}0\\ 0\\ 1\\ \end{array}\right).$ (10)
The necessary assumptions i) and ii), cf. with section II-B, are fulfilled since the state-space realization (10) is both fully controllable and observable, and one can guarantee that the system matrix $A$ is Hurwitz under the condition that inequality (4) is satisfied. The feedback nonlinearity
$\phi(z,t)=c\,d\,\exp\bigl{(}e|x_{2}|\bigr{)}\,z$ (11)
proves to satisfy the necessary assumption iii), since it belongs to the sector $[0,k]$ with $k=cd>0$ for all $|x_{2}(t)|\in\mathbb{R}_{+}$. Fulfillment of the necessary assumption iv) follows from the fact that (11) is single-valued with respect to the $(z,|x_{2}|)$ pair, i.e. it is memoryless, and continuous in $t$ and Lipschitz in $z$. Two of the last statements are due to the exponential term, and since $x_{2}(t)\in\mathcal{C}^{3}$ for all $u(t)$ and $\sigma(t)$ to be Lipschitz. Now, we are in position to show the absolute stability of the unperturbed control system (1), (8).
Theorem 2
Given an unperturbed feedback system (1), (8). Then, its origin is globally asymptotically stable if
$\frac{c^{2}d-acd\,\omega^{2}}{(b\omega-\omega^{3})^{2}+(c-a\,\omega^{2})^{2}}>-1,\quad\forall\>\omega.$ (12)
Proof:
Substituting (10) into (6) and evaluating the loop transfer function for $s=j\omega$ results in
$H(j\omega)=-\frac{1}{a\,\omega^{2}-c+j\omega\,(\omega^{2}-b)}.$ (13)
Evaluating the real part of (13) and substituting it, together with $k=cd$, into (7) results in
$01+\frac{c^{2}d-acd\,\omega^{2}}{(b\omega-\omega^{3})^{2}+(c-a\,\omega^{2})^{2}}>$ 0 .
Comparing it to (12), this completes the proof. ∎
Remark 1
For both boundary frequencies $\omega=0$ and $\omega=\infty$ the condition (12) is always fulfilled. Indeed, for $\omega=\infty$ one obtains $-1/\infty>-1$. Also, substituting $\omega=0$ into (12) results in $d>-1$, that is always true for the control (8).
Figure 1: Visualization of the circle condition for $H(j\omega)$ and sector $[0,cd]$.
A graphical visualization of the circle stability criterion applied to $H(j\omega)$, with an exemplary stable characteristic polynomial (3) satisfying (12) and $k=cd$, is depicted in Fig. 1 for the sake of explainability.
III-B Disturbance rejection
Consider a standard feedback control loop as shown in Fig. 2. Note that if a reference $r(t)$ is also applied, then both controllers (2) and (8) will incorporate the control error $\epsilon=r-y$ instead of $y$ and, respectively, its derivative and integral. Consequently, they will change to additive inverse, i.e. $-a\dot{y}$ changes to $a\dot{\epsilon}$ and so on. In the following, for the sake of simplicity and due to the focus on disturbance rejection, we assume $r=0$. Same time, $\sigma\neq 0$ is entering the loop. Further, the control (8) is denoted by nl-PID while its linear counterpart (2) by PID, respectively, cf. Fig. 2.
Figure 2: Closed loop system with (nl-)PID control.
While the system plant transfer function is given by
$G(s)=\frac{1}{s^{2}},$ (14)
one can temporary assume (9) as a ’frozen’ (i.e. constant) parameter and write the control transfer function as
[…]
The corresponding perturbation-to-output transfer function is
$S_{\sigma y}(s)=\frac{y(s)}{\sigma(s)}=\frac{G(s)}{1+C(s)G(s)}.$ (16)
Substituting (14), (15) and evaluating (16), one can recognize that for any $\sigma=\mathrm{const}$, the resulting $y(s)\rightarrow 0$ as $s\rightarrow 0$, i.e. at steady-state. This is irrespectively of the $\Omega$-value for which the characteristic polynomial (3) with $c=\Omega$ is guaranteed to remain stable. Evaluating the magnitude of the frequency response function $S_{\sigma y}(j\omega)$ one obtains
$\bigl{|}S_{\sigma y}(j\omega)\bigr{|}=\frac{\omega}{\sqrt{\bigl{(}b\omega-\omega^{3}\bigr{)}^{2}+\bigl{(}\Omega-a\omega^{2}\bigr{)}^{2}}}\,.$ (17)
One can show that for an increasing $\Omega$, while satisfying $\Omega<ab$, the peak value of (17) also increases, i.e. $\Omega\nearrow\>\Rightarrow\>\max\,\bigl{|}S_{\sigma y}(j\omega)\bigr{|}\nearrow$. A qualitative behavior of (17) is shown in Fig. 3 for the exemplary assigned $\Omega\in\{1c,2c,3c,4c\}$.
Figure 3: Magnitude of frequency response function $S_{\sigma y}$ depending on $\Omega$.
Same time, one can also show that for frequencies lower than the peak frequency $\omega_{\mathrm{peak}}$, an increasing $\Omega$ leads to a decreasing magnitude value, i.e. $\Omega\nearrow\>\Rightarrow\>\bigl{|}S_{\sigma y}(j\omega)\bigr{|}\searrow$ for $\omega\ll\omega_{\mathrm{peak}}$, cf. Fig. 3. Here we recall that higher peak values give rise to higher overshoots during the transients, like for example in case when all frequencies are excited by a stepwise $\sigma$. On the contrary, lower $|S_{\sigma y}|$ values reduce the ultimately bounded output error, correspondingly, improve the convergence performance. Thus, an appropriately varying $\Omega$ value, cf. (9), can improve both the transient and steady-state behavior of the control system. We should recall, however, that the behavior (17), cf. also Fig. 3, discloses only a steady-state response of the feedback control system.
In order to assess the transient response of the control system to a matched perturbation $\sigma$, and that depending on the $\Omega$ parameter, we apply the unit step $\sigma(s)$ so that
$y(s)=\frac{G(s)}{1+C(s)G(s)}\cdot\frac{1}{s}.$ (18)
In the following, for the sake of simplicity when developing and analyzing the time domain solution of (18) we assume the real eigenvalues $\lambda_{1},\lambda_{2}>0$ so that the stable transfer function (16) has a double-pole pair at $-\lambda_{1}$ and, respectively, one pole at $-\lambda_{2}$. One can show that such poles configuration corresponds to the control gains equal to
$a=2\lambda_{1}+\lambda_{2},\quad b=\lambda_{1}^{2}+2\lambda_{1}\lambda_{2},\quad\Omega=\lambda_{1}^{2}\lambda_{2}.$ (19)
Solving (18) in time domain and putting outside the brackets the common term $\Gamma=(\lambda_{1}-\lambda_{2})^{-2}$ results in
$y(t)=\Gamma\Bigl{(}\exp(-\lambda_{2}t)-\bigl{(}1-(\lambda_{2}-\lambda_{1})\,t\bigr{)}\exp(-\lambda_{1}t)\Bigr{)}.$ (20)
Expressing $\lambda_{2}$ in terms of $\Omega$, cf. (19), one obtains from (20)
$y(t)$ $=$ $\hat{\Gamma}\biggl{(}\exp\Bigl{(}-\frac{\Omega}{\lambda_{1}^{2}}t\Bigr{)}-\Bigl{(}1-\frac{\Omega-\lambda_{1}^{3}}{\lambda_{1}^{2}}\,t\Bigr{)}\exp\bigl{(}-\lambda_{1}t\bigr{)}\biggr{)},$
$\hat{\Gamma}$ $=$ $\frac{\lambda_{1}^{4}}{(\lambda_{1}^{3}-\Omega)^{2}}.$ (21)
Analyzing (21), one can show that an increasing $\Omega$ (for a fixed $\lambda_{1}$ value) leads to a decreasing peak in the transient response of $y(t)$, cf. Fig. 4 for exemplary $\Omega\in\{0.5c,1c,2c\}$.
Figure 4: Exemplary transient response of (21) for varying $\Omega$.
Numerical example
The following numerical example visualizes convergence of the control system (1), (8) for various parameter settings. The assigned (most simple) disturbance is a constant value $\sigma=-100$. The set linear control parameters are $a=60$, $b=1100$, $c=3000$, resulting in a poles configuration $\lambda\in\{-3.28,\,-28.36\pm j10.47\}$, i.e. when the nonlinear integral gain is excluded, i.e. $d=0$. Note that for an integral gain equal to $60002c=$ 6000, that corresponds to $d=1$ when $y\rightarrow 0$, the poles configuration results in $\lambda^{*}\in\{-10,\,-20,\ -30\}$. The initial value $y(0)=-1$ is assigned for a simultaneous set-point and disturbance rejection control task.
For a fixed parameter $e=-10$, the output convergence on the logarithmic scale is compared for the set parameter values $d=\{1,2,3\}$ in Fig. 5. The linear PID case, i.e. for $d=0$, is also included for the sake of comparison.
Figure 5: Output convergence on logarithmic scale for varying $d$.
Next, for a fixed parameter $d=2$, the output convergence on the logarithmic scale is compared for the set parameter values $e=\{-10,-100,-1000\}$ in Fig. 6. Also here, the linear PID case of $d=0$ is shown for comparison.
Figure 6: Output convergence on logarithmic scale for varying $e$.
One can recognize that for some lower threshold $\varepsilon=\mathrm{const}$ of the residual error, the convergence of the nl-PID control outperforms that one of the PID, in terms of $|y(t)|<\varepsilon$ for $t>t_{0}$ with $t_{0}$ to be a metric of the settling time. Simultaneously, for higher $\varepsilon$ values, for instance with $t_{0}<1$ sec cf. Figs. 5, 6, it is a combination of $(d,e)$ parameters that allows shaping the output convergence to be superior comparing to the linear case. Note that in addition to $y(t)$ convergence, the transient over- or undershoots and the control effort can appear as essential in a particular application.
III-C Extension by an actuator dynamics
When the second-order system plant (1) is extended by a (parasitic) actuator dynamics
$v(s)=F(s)u(s),$ (22)
where $F(s)$ is a stable and strictly proper transfer function with $F(0)\rightarrow\mathrm{const}$, the system (1) expands to
$\ddot{y}(t)=v(t)+\sigma(t).$ (23)
Then, the extended plant (23), (22) is controlled in the same way by (8), provided the resulted Lur’e system equally fulfills the necessary assumptions i)-iv), cf. with section II-B.
For the extended plant (23), (22), the corresponding input-output loop transfer function results then in
$H(s)=\frac{F(s)}{s^{3}+as^{2}+bs+c}.$ (24)
Since neither the structure of a feedback system nor the nonlinearity (11) change by (22), the same circle criterion as given in section II-B is applicable to (24). Respectively, the parametric inequality (12) of Theorem 2 should modify, so that to take into account also the polynomial coefficients of $F(j\omega)$. Since both the absolute stability and disturbance rejection for (24) follow the same lines of argumentation as in sections III-A, III-B, and a particular actuator dynamics $F(s)$ is rather application specific, we omit a detailed development here. At the same time, it is worst emphasizing that the experimental case study provided below in section IV has the second-order system plant additionally affected by such first-order actuator dynamics $v(s)=\kappa(s\mu+1)^{-1}u(s)$, with a gain factor $\kappa$ and (not negligible) time constant $\mu>0$.
IV EXPERIMENTAL EVALUATION
The proposed extended nonlinear PID control given by (8) is experimentally evaluated and additionally benchmarked with a standard PID control (2). In addition, the corresponding PD control (i.e. $c=0$) is also evaluated for highlighting the level of perturbations in the system at hand and, this way, disclosing necessity of the integral control action. The common control parameters are assumed the same, while all parameter values are summarized in Table I.
TABLE I: Control parameters
control \  params. $a$ $b$ $c$ $d$ $e$
PD $10$ $1000$
PID $10$ $1000$ $5000$
nl-PID $10$ $1000$ $5000$ $1$ $-100$
The experimental laboratory setup used for evaluation of all three controllers is depicted in Fig. 7.
Figure 7: Experimental setup of 1DOF actuator system (laboratory view).
The motion system consists of a voice-coil driven actuator with 1DOF and contactless measured output position $y(t)$. The latter contains a relatively hight level of the sensing noise. The robust finite-time second-order sliding-mode differentiator [25, 26] is used for obtaining $\dot{y}(t)$. The sampling frequency of the implemented real-time control is set to 10 kHz. A more detailed description of the experimental system, including the identified system parameters, can be found in [18].
Further it is worth noting that: i) an additional first-order actuator dynamics (due to electro-magnetic behavior of the voice-coil motor) is in place, and ii) the unknown perturbation $\sigma(t)$ is mainly due to the gravity, nonlinear friction, and force ripples of the anchor-stator interaction.
The measured output response to the step reference $r=0.008$ m, that is applied at time $t=1$ sec, is shown in Fig. 8 for the nl-PID, PID, and PD controllers.
Figure 8: Measured output response of the nl-PID, PID, and PD controls.
The absolute output error $|\epsilon|\equiv|r-y|$ of the nl-PID and PID controllers is additionally visualized on the logarithmic scale in Fig. 9. One can recognize that for $|\epsilon|<0.0001$ m, both controllers have to deal with a sensing noise so that the residual control errors are similar. At the same time, the convergence performance is well distinguishable for the time range between 1.18 and 1.7 sec. The transient behavior (including over- and undershoots) at the time between 1 and 1.18 sec is well comparable.
Figure 9: Absolute output error (logarithmic) of the nl-PID, PID controls.
V CONCLUSIONS
In this paper, a novel nonlinear extension of the PID feedback control (denoted by nl-PID) is introduced. The nonlinear extension is proposed for integral control action that allows improving convergence performance in presence of the matched unknown perturbations in the second-order systems. The nonlinear integral control part does not change the system structure and is continuous in time and Lipschitz in the system output variable, so that the absolute stability by means of the circle criterion turned out applicable and sufficient for analysis. The nl-PID control has five design parameters in total, while keeping the same meaning and tuning rules for the proportional, integral, and derivative feedback gains. The global asymptotic stability of the nl-PID control was shown for the case of constant perturbations. For the case of Lipschitz perturbations, an ultimately bounded output error can be guaranteed by consideration in frequency domain. Beyond the analysis and numerical examples, an experimental case study is demonstrated for benchmarking the standard PD, PID, and proposed nl-PID controls.